OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

\[y=\left(\begin{matrix}2 \\ 3\end{matrix}\right)x+400\]

OpenStudy (anonymous):

Thank you then would the slope be 2/3 and the y intercept would be (0,400) ?

OpenStudy (anonymous):

@johnweldon1993 is this right?

OpenStudy (anonymous):

slope is -2/3 and the y intercept is as you said (0,400) then I would use the slope to do the rest or go 2 points to the right and 3 to the down. After this it should look like 3,398

OpenStudy (anonymous):

the last part confused me /.\

OpenStudy (johnweldon1993):

Right^ To change \(\large 2x + 3y = 1200\) into slope intercept form...subtract 2x from both sides \[\large 3y = -2x + 1200\] divide by 3 \[\large y = -\frac{2}{3} + 400\]

OpenStudy (johnweldon1993):

Argh forgot the 'x' -_- \[\large y = -\frac{2}{3}x + 400\] There we go :)

OpenStudy (anonymous):

thanks guys (:

OpenStudy (anonymous):

Anytime!

OpenStudy (johnweldon1993):

Of course!